A steel rod is 3.418 cm in diameter at 26.00°C. A brass ring has an interior diameter of 3.410 cm at 26.00°C. At what common temperature will the ring just slide onto the rod? The linear expansion coefficient of steel is 11.00 × 10-6 1/C°. The linear expansion coefficient of brass is 19.00 × 10-6 1/C°.
Answer #1

The ring will just slide onto the rod when diameters of both become same.

Let at temperature T diameters of both are same.

let Ds and DB are the initial diameters of steel rod and brass ring respectively.While Ds' and DB' are the diameters of Steel rod and brass ring at temperature T respectively.

we can write

Ds= Ds (1as (T 26)) --------(1)

DR DB(1a (T 26))   ---------(2)

where as and \alpha_{B} are the linear expansion coefficient for steel and brass respectively.

According to question-

Ds DR B S.

\RightarrowDs(1as (T-26)) DB(1 a (T- 26)

\Rightarrow3.418 x (111 x 10(T- 26)) 3.410(1 19 x 10(T- 26))

\Rightarrow3.418 x 11 x 105(T - 26) 3.410 x 19 x 10(T - 26)) = 3.410 3.418

\Rightarrow27.192 x 10- (T 26) -0.008

T-26 294.20T=268.20^{0}C

ET = 320.20°C